A flower bed measuring 8 feet by 10 feet is bordered by a concrete walk that is 2 feet wide. What is the are of the concrete walk?
Answer by nerdybill(7384)

A flower bed measuring 8 feet by 10 feet is bordered by a concrete walk that is 2 feet wide. What is the are of the concrete walk?
.
area of flower bed: 8*10 = 80 sq ft
area of flower bed and walk: (8+4)(10+4)=12*14=168 sq ft
area of walk = "area of flower bed and walk" - "area of flower bed"
area of walk = 168 - 80 = 88 sq ft